OpenStudy (anonymous):

f(x) = tan (pi times x)/2. Find the x-values at which f is not continuous. Use k as an arbitrary integer.

OpenStudy (nikvist):

\[f(x)=\tan{\frac{\pi x}{2}}\quad or\quad f(x)=\frac{\tan{\pi x}}{2}\quad ???\]

OpenStudy (anonymous):

The first one.